# No. 3-11, Wenhua Road, Shenyang, 110819, China

### 03 Mar 2024 v1

@@ 1 Introduction
The PBH abundance $f_{\rm PBH}$ can be calculated from the power spectrum ${\cal P}_{\cal R}(k)$ of the primordial curvature perturbation ${\cal R}$ PS ; peak ; Green:2004wb . On large scales (e.g., the pivot scale $k_{*}=0.05~{}{\rm Mpc}^{-1}$ in the Planck satellite experiment), ${\cal P}_{\cal R}$ is precisely measured by the anisotropies of the cosmic microwave background (CMB), with an amplitude of $2.10\times 10^{-9}$ Planck . However, to produce abundant PBHs, ${\cal R}$ must be large enough, so that ${\cal P}_{\cal R}$ is significantly enhanced up to ${\cal O}(10^{-2})$ on small scales. In the usual single-field slow-roll (SR) inflation models, such a huge enhancement of ${\cal P}_{\cal R}$ is almost impossible. Therefore, the SR conditions must be violated on small scales, and various ultraslow-roll (USR) inflation models are thus proposed Garcia-Bellido:2017mdw ; Kannike:2017bxn ; Germani:2017bcs ; Motohashi:2017kbs ; Dimopoulos:2017ged ; Ezquiaga:2017fvi ; Ballesteros:2017fsr ; Cicoli:2018asa ; Ozsoy:2018flq ; Dalianis:2018frf ; Ballesteros:2018wlw ; Cheng:2018qof ; Bhaumik:2019tvl ; Mahbub:2019uhl ; liu ; Mishra:2019pzq ; Cai:2019bmk ; Figueroa:2020jkf ; Ragavendra:2020sop ; Choi:2021yxz ; Kefala:2020xsx ; Lyc ; Wq ; Dalianis:2021iig ; Inomata:2021uqj ; Cheng:2021lif ; Wu:2021mwy ; Inomata:2021tpx ; Figueroa:2021zah ; Hooshangi:2022lao ; Geller:2022nkr ; Franciolini:2022pav ; Balaji:2022zur ; Raveendran:2022dtb ; Bhatt:2022mmn ; Gu:2022pbo ; Zjx ; Mu:2022dku ; Ragavendra:2023ret . In these models, there is always a plateau or a (near-) inflection point on the inflaton potential, where the inflaton rolls down extremely slowly, amplifying ${\cal P}_{\cal R}$ accordingly.
Albeit the amplitude and shape of ${\cal P}_{\cal R}$ can be straightforwardly obtained in various USR inflation models by numerical methods, the physical comprehension behind seems still indistinct. Previously, in Refs. Carrilho:2019oqg ; liu ; Zhai:2023azx , the spectral index of the steepest growth of ${\cal P}_{\cal R}$ was analytically investigated, and in Ref. Franciolini:2023lgy , the authors raised an approximation method to calculate the Fourier mode ${\cal R}_{k}$ of the primordial curvature perturbation. However, to our knowledge, these analyses are still not sufficient in some aspects. For example, there is always a sharp dip in ${\cal P}_{\cal R}$ before its steep growth. In some early literature Byrnes:2018txb ; Ballesteros:2020sre ; Tasinato:2020vdk ; Ozsoy:2021qrg ; Ozsoy:2021pws ; Cole:2022xqc ; Karam:2022nym ; Pi:2022zxs ; shuzhi , the location of the dip was calculated numerically, but the underlying physical explanation is lacking yet. Although the detail of this dip does not affect the PBH abundance or SIGW spectrum, as they are relevant mainly to the peak of ${\cal P}_{\cal R}$, we still hope to answer whether the minimum of ${\cal P}_{\cal R}$ is zero analytically.

### 10 Mar 2024 v2

@@ 1 Introduction
The PBH abundance $f_{\rm PBH}$ can be calculated from the power spectrum ${\cal P}_{\cal R}(k)$ of the primordial curvature perturbation ${\cal R}$ PS ; peak ; Green:2004wb . On large scales (e.g., the pivot scale $k_{*}=0.05~{}{\rm Mpc}^{-1}$ in the Planck satellite experiment), ${\cal P}_{\cal R}$ is precisely measured by the anisotropies of the cosmic microwave background (CMB), with an amplitude of $2.10\times 10^{-9}$ Planck . However, to produce abundant PBHs, ${\cal R}$ must be large enough, so that ${\cal P}_{\cal R}$ is significantly enhanced up to ${\cal O}(10^{-2})$ on small scales. In the usual single-field slow-roll (SR) inflation models, such a huge enhancement of ${\cal P}_{\cal R}$ is almost impossible. Therefore, the SR conditions must be violated on small scales, and various ultraslow-roll (USR) inflation models are thus proposed Garcia-Bellido:2017mdw ; Kannike:2017bxn ; Germani:2017bcs ; Motohashi:2017kbs ; Dimopoulos:2017ged ; Ezquiaga:2017fvi ; Ballesteros:2017fsr ; Cicoli:2018asa ; Ozsoy:2018flq ; Dalianis:2018frf ; Ballesteros:2018wlw ; Cheng:2018qof ; Bhaumik:2019tvl ; Mahbub:2019uhl ; liu ; Mishra:2019pzq ; Cai:2019bmk ; Figueroa:2020jkf ; Ragavendra:2020sop ; Choi:2021yxz ; Kefala:2020xsx ; Lyc ; Wq ; Dalianis:2021iig ; Inomata:2021uqj ; Cheng:2021lif ; Wu:2021mwy ; Inomata:2021tpx ; Figueroa:2021zah ; Hooshangi:2022lao ; Geller:2022nkr ; Franciolini:2022pav ; Balaji:2022zur ; Raveendran:2022dtb ; Bhatt:2022mmn ; Gu:2022pbo ; Zjx ; Mu:2022dku ; Ragavendra:2023ret ; Balaji:2022rsy ; Qin:2023lgo . In these models, there is always a plateau or a (near-) inflection point on the inflaton potential, where the inflaton rolls down extremely slowly, amplifying ${\cal P}_{\cal R}$ accordingly.
Albeit the amplitude and shape of ${\cal P}_{\cal R}$ can be straightforwardly obtained in various USR inflation models by numerical methods, the physical comprehension behind seems still indistinct. Previously, in Refs. Carrilho:2019oqg ; liu ; Zhai:2023azx , the spectral index of the steepest growth of ${\cal P}_{\cal R}$ was analytically investigated, and in Ref. Franciolini:2023lgy , the authors raised an approximation method to calculate the Fourier mode ${\cal R}_{k}$ of the primordial curvature perturbation. However, to our knowledge, these analyses are still not sufficient in some aspects. For example, there is always a sharp dip in ${\cal P}_{\cal R}$ before its steep growth. In some early literature Byrnes:2018txb ; Ballesteros:2020sre ; Tasinato:2020vdk ; Ozsoy:2021qrg ; Ozsoy:2021pws ; Cole:2022xqc ; Karam:2022nym ; Pi:2022zxs ; shuzhi ; Domenech:2023dxx ; Alho:2020cdg , the location of the dip was calculated numerically, but the underlying physical explanation is lacking yet. Although the detail of this dip does not affect the PBH abundance or SIGW spectrum, as they are relevant mainly to the peak of ${\cal P}_{\cal R}$, we still hope to answer whether the minimum of ${\cal P}_{\cal R}$ is zero analytically.